OpenStudy (anonymous):

will someone please give me a basic idea on how to do a question like 5w+6=9-w please use words so i can understand it better, thanks.

OpenStudy (anonymous):

are trying to find out what w is?

OpenStudy (anonymous):

Well what exactly are you having difficulty with when you first look at a problem like this?

OpenStudy (anonymous):

Well when solving things like this, you're trying to get w alone so you can find its value. By doing this, you have to undo things. So, you do opposite operations to get w. So with this, first you have to get w on the same side, and then you can start finding w. So: 5w+6=9-w First, add w to both sides. (see how it's -w? always take the sign that's in front.) So: 5w+6=9-w +w +w This cancels the w on the right side. 6w+6=9 Now, we need to get the +6 away, by doing so we do the opposite, which is -6. Remeber, whatever you do to one side, you MUST do to the other. 6w+6=9 -6 -6 6w=3 Now, 6w is multiplying, so the opposite would be to divide. So, to get w alone, we'll divide both side by 6. 6w=3 /6 /6 w=0.5 You can put this into your original equation to check it: 5w+6=9-w 5(0.5)+6=9-0.5 2.5+6=8.5 8.5=8.5 Hope that makes sense and helps you. :o)